a) Two identical charges q1 and q2 make up a system with a potential energy of 1590J when the distance between them is 0.4m. Find the magnitude of the charges q1 and q2.



1
Expert's answer
2021-10-18T11:13:17-0400
"U=k\frac{q^2}{r}\\1590=9\cdot10^{9}\frac{q^2}{0.4}\\q=2.66\cdot10^{-4}\ C"